### v1

@@ Abstract
The recognition of art styles and authors is crucial in areas like cultural heritage protection, art market analysis, and historical research. With the advancement of deep learning, Convolutional Neural Networks (CNNs) and Transformer models have become key tools for image classification. While CNNs excel in local feature extraction, they struggle with global context, and Transformers are strong in capturing global dependencies but weak in fine-grained local details. To address these challenges, this paper proposes a fusion model combining CNNs and Transformers for art style and author recognition. The model first extracts local features using CNNs, then captures global context with a Transformer, followed by a feature fusion mechanism to enhance classification accuracy. Experiments on Chinese and oil painting datasets show the fusion model outperforms individual CNN and Transformer models, improving classification accuracy by 9.7% and 7.1%, respectively, and increasing F1 scores by 0.06 and 0.05. The results demonstrate the model’s effectiveness and potential for future improvements, such as multimodal integration and architecture optimization.
@@ 1 Introduction
Art recognition, particularly in terms of style and author classification, is a critical research task in computer vision, with widespread applications in cultural heritage preservation, art market regulation, and historical research. With the rapid development of the internet, the digitization of artworks has enabled a large number of art pieces to be displayed and traded online. Consequently, automatic classification and recognition of these artworks in terms of style and author have become key issues of interest in both industry and academia.
Style and author recognition not only aids museums and galleries in accurately categorizing artworks but also provides technical support for the regulated management of the art market. With intelligent recognition systems, users can efficiently filter artworks that meet their needs from a large collection. Moreover, the classification of art styles and authors offers new perspectives for art scholarship, especially in areas like the authentication of artworks and the identification of creative periods, which have significant practical implications.
In recent years, deep learning techniques have made remarkable progress in image classification tasks, and Convolutional Neural Networks (CNNs) have become the mainstream approach for art style and author classification. CNNs automatically extract meaningful features from images through multi-layer convolution operations, overcoming the limitations and instability faced by traditional hand-crafted feature extraction methods (e.g., SIFT, HOG). The emergence of deeper networks, such as ResNet and Inception, has significantly improved image recognition accuracy.
@@
    The model is tested on datasets from both Chinese painting and oil painting domains, enabling it to handle diverse art styles and author classification tasks.
@@ 2 Related Work
2.1 Art Recognition
Art recognition, particularly for identifying artistic styles and authors, has a long history, initially relying on traditional image processing techniques. Early methods focused on handcrafted features, including color histograms, texture descriptors such as Gabor filters, and shape features like Histogram of Oriented Gradients (HOG) and Scale-Invariant Feature Transform (SIFT). These approaches were foundational in the field but have limitations when applied to complex artworks, where subtle variations and intricate stylistic elements are common.
@@ 2.1 Art Recognition
The advent of deep learning techniques, particularly Convolutional Neural Networks (CNNs), has significantly advanced the field of art recognition[2]. CNNs, unlike traditional methods, can automatically learn hierarchical features from images, enabling them to capture complex patterns, textures, and relationships in a way that manual feature extraction cannot. In art recognition, CNNs have shown significant promise. For instance, Dosovitskiy et al. (2016) proposed a method for classifying art styles using CNNs, achieving excellent results in style recognition tasks. CNNs, such as VGG (Simonyan & Zisserman, 2014)[3] and ResNet (He et al., 2016)[4], have demonstrated their ability to automatically learn features that are highly useful for distinguishing between art styles and authors. However, these methods typically extract local features and lack the ability to model global information, which is essential for understanding complex artworks that contain multiple interconnected elements, colors, and textures.
@@ 2.2 Application of Transformer in Visual Tasks
Despite the success of Transformers in many image classification tasks, they have notable challenges. ViT models, for instance, typically require large datasets for effective training, as they do not perform well on small datasets compared to CNNs. This is primarily due to the large number of parameters in Transformers, which require extensive training data to generalize effectively. Additionally, Transformers, while excellent at capturing global information, often struggle with local feature extraction, which is crucial for tasks such as fine-grained art style recognition.
@@ 2.3 Related Research on CNN-Transformer Fusion
One prominent approach is proposed by Xie et al., who introduced a hybrid model [7]that first uses CNNs to extract low-level features from images and then passes these features to a Transformer for global modeling. This model has demonstrated good performance in various image classification tasks, including facial recognition and general object classification. By combining CNNs’ local feature extraction with Transformers’ ability to capture long-range dependencies, the model achieves enhanced performance over using either architecture in isolation.
Liu et al. (2021) also explored the fusion of CNNs and Transformers, focusing on a multi-scale feature fusion approach. In this method, CNNs extract features at different scales, and these features are subsequently fused with the global features captured by the Transformer. This multi-scale approach is particularly useful when dealing with small datasets, as it allows the model to capture both local and global features from images at different resolutions. This fusion technique has shown significant improvements in small-data environments, making it highly relevant for art recognition tasks, where high-quality labeled datasets may be scarce.
While the fusion of CNNs and Transformers has shown promising results in various applications, the integration of these models for the specific task of art style and author classification remains an area in need of further exploration. Existing research on hybrid models primarily focuses on generic image classification tasks, and there has been limited investigation into optimizing CNN-Transformer fusion architectures for the unique challenges posed by artwork recognition, such as capturing fine brushstrokes, identifying complex color gradients, and distinguishing between different artistic styles and authors. Therefore, further work is needed to refine the fusion strategies and adapt them to the specific nuances of art recognition tasks.
@@ 3.2 Model Architecture
The model proposed in this study is built on a hybrid architecture combining Convolutional Neural Networks (CNNs) and Transformers, designed to leverage the complementary strengths of these two deep learning approaches. The hybrid model aims to improve classification accuracy and robustness, particularly for the nuanced and diverse nature of art recognition tasks.
@@
Fusion Strategy: The hybrid architecture employs a cascading fusion approach to combine CNN and Transformer components. After the CNN extracts local features, these features are passed to the Transformer for further processing. The Transformer applies multiple layers of self-attention to integrate and enhance the global dependencies between these local features. Finally, the local and global features are concatenated and passed to a fully connected layer for classification. This fusion strategy allows the model to effectively combine the strengths of both CNNs and Transformers—local feature extraction and global context modeling—resulting in better performance in classifying art styles and authors.
@@ 3.3 Loss Function and Optimization
Loss Function: The model employs the cross-entropy loss function, which is well-suited for multi-class classification tasks. Given the nature of the art recognition task, the model outputs a probability distribution for each sample, representing the likelihood of it belonging to a particular style or author. The cross-entropy loss function calculates the difference between the predicted probability distribution and the true label distribution, guiding the model to minimize this discrepancy during training. To address the class imbalance in the dataset, where some classes (e.g., rare styles or authors) may have fewer samples, we use a weighted cross-entropy loss. The weights are calculated based on the inverse frequency of each class in the dataset, ensuring that the model gives appropriate attention to minority classes.
@@ 3.4 Data Augmentation and Regularization
Regularization: To further combat overfitting and improve the model’s generalization to unseen data, we employ dropout regularization during training, with a keep probability of 0.5. This technique randomly deactivates certain neurons during training, preventing the model from becoming overly reliant on any single feature and promoting more robust learning. Additionally, we apply L2 regularization (also known as weight decay) to the fully connected layers, with a penalty coefficient of 0.01. This further discourages overly large weights, helping the model maintain generalization across different art styles and authors.
@@ 4.3 Experimental Results
The experimental results on both the Chinese and oil painting datasets are summarized in the table 1. From these results, it is clear that the fusion model significantly outperforms both the individual CNN and Transformer models across all evaluation metrics. For the Chinese painting dataset, the fusion model achieved an impressive 9.7% increase in accuracy and a 0.09 increase in F1-score compared to the CNN-only model. Similarly, for the oil painting dataset, the fusion model outperformed the CNN and Transformer models by 7.1% in accuracy and 0.05 in F1-score. These results highlight the importance of combining the strengths of both CNNs and Transformers for art recognition tasks, where both local and global features are essential for accurate classification.
@@ 4.4 Experimental Analysis
Fusion Model (CNN + Transformer): The fusion model, which combines the local feature extraction capability of the CNN with the global modeling power of the Transformer, achieves the best overall performance on both datasets. By integrating the strengths of both architectures, the fusion model captures both local details and global context effectively. This allows the model to handle the nuances of different art styles and authors more effectively. The significant improvements in accuracy and F1-score demonstrate the effectiveness of this approach, particularly when dealing with complex art forms like Chinese and oil paintings.
Furthermore, the fusion model exhibits improved robustness, especially in small-data environments. The integration of both local and global features helps the model generalize better, even when fewer training examples are available, which is particularly important in art recognition tasks where labeled data can be limited.
@@ 4.5 Discussion
The fusion of CNN and Transformer allows the model to fully leverage the strengths of both. CNNs are adept at extracting local details from images, such as brushstrokes and color distribution, while Transformers can capture global dependencies and understand the overall artistic style and author intent. Through the cascading fusion strategy, the model can focus on both local details and global context, thereby improving classification accuracy.
@@ 5 Conclusion
This study proposes an innovative fusion model that combines Convolutional Neural Networks (CNN) and Transformers to address the challenges of art style and author recognition. CNNs, known for their exceptional ability to capture local features such as textures, brushstrokes, and color distributions, are used as the initial step to process the artwork images. Transformers, on the other hand, excel at modeling long-range dependencies and global relationships, which are crucial for understanding the overall composition and style of an artwork. By integrating these two powerful models, the proposed fusion approach leverages the strengths of both CNNs and Transformers, allowing the model to effectively extract detailed local features while also capturing the broader contextual information.
@@
Additionally, the proposed model is especially effective in small-data environments, where the ability to generalize from limited data is crucial. The fusion model not only provides a significant boost in performance but also shows potential in enhancing the recognition of subtle artistic nuances that are often challenging for standalone models.
Looking ahead, future research will focus on optimizing the model architecture to further improve its accuracy and efficiency. This could involve exploring more advanced convolutional and Transformer architectures, such as deeper CNN networks or novel Transformer variants. Additionally, the incorporation of multimodal information, including textual descriptions or historical data about artists, could further enrich the model’s understanding of artworks, providing a more holistic approach to art recognition. Expanding the dataset to include a wider variety of art styles and authors will also be a key area for future work, helping to improve the model’s generalizability and adaptability across different domains of art.

### v2

@@ Abstract
The identification of art authors is crucial in areas like cultural heritage protection, art market analysis, and historical research. With the advancement of deep learning, Convolutional Neural Networks (CNNs) and Transformer models have become key tools for image classification. While CNNs excel in local feature extraction, they struggle with global context, and Transformers are strong in capturing global dependencies but weak in fine-grained local details. To address these challenges, this paper proposes a fusion model combining CNNs and Transformers for art author identification. The model first extracts local features using CNNs, then captures global context with a Transformer, followed by a feature fusion mechanism to enhance classification accuracy. Experiments on Chinese and oil painting datasets show the fusion model outperforms individual CNN and Transformer models, improving classification accuracy by 9.7% and 7.1%, respectively, and increasing F1 scores by 0.06 and 0.05. The results demonstrate the model’s effectiveness and potential for future improvements, such as multimodal integration and architecture optimization.
@@ 1 Introduction
Art identification, particularly in terms of author classification, is a critical research task in computer vision, with widespread applications in cultural heritage preservation, art market regulation, and historical research. With the rapid development of the internet, the digitization of artworks has enabled a large number of art pieces to be displayed and traded online. Consequently, automatic classification and identification of these artworks in terms of author have become key issues of interest in both industry and academia.
Author identification not only aids museums and galleries in accurately categorizing artworks but also provides technical support for the regulated management of the art market. With intelligent identification systems, users can efficiently filter artworks that meet their needs from a large collection. Moreover, the classification of art authors offers new perspectives for art scholarship, especially in areas like the authentication of artworks and the identification of creative periods, which have significant practical implications.
In recent years, deep learning techniques have made remarkable progress in image classification tasks, and Convolutional Neural Networks (CNNs) have become the mainstream approach for art author classification. CNNs automatically extract meaningful features from images through multi-layer convolution operations, overcoming the limitations and instability faced by traditional hand-crafted feature extraction methods (e.g., SIFT, HOG). The emergence of deeper networks, such as ResNet and Inception, has significantly improved image identification accuracy.
@@
    The model is tested on datasets from both Chinese painting and oil painting domains, enabling it to handle diverse art author classification tasks.
@@ 2 Related Work
2.1 Art Identification
Art identification, particularly for identifying artistic authors, has a long history, initially relying on traditional image processing techniques. Early methods focused on handcrafted features, including color histograms, texture descriptors such as Gabor filters, and shape features like Histogram of Oriented Gradients (HOG) and Scale-Invariant Feature Transform (SIFT). These approaches were foundational in the field but have limitations when applied to complex artworks, where subtle variations and intricate stylistic elements are common.
@@ 2.1 Art Identification
The advent of deep learning techniques, particularly Convolutional Neural Networks (CNNs), has significantly advanced the field of art identification[2]. CNNs, unlike traditional methods, can automatically learn hierarchical features from images, enabling them to capture complex patterns, textures, and relationships in a way that manual feature extraction cannot. In art identification, CNNs have shown significant promise. For instance, Dosovitskiy et al. (2016) proposed a method for classifying art styles using CNNs, achieving excellent results in style identification tasks. CNNs, such as VGG (Simonyan & Zisserman, 2014)[3] and ResNet (He et al., 2016)[4], have demonstrated their ability to automatically learn features that are highly useful for distinguishing between art authors. However, these methods typically extract local features and lack the ability to model global information, which is essential for understanding complex artworks that contain multiple interconnected elements, colors, and textures.
@@ 2.2 Application of Transformer in Visual Tasks
Despite the success of Transformers in many image classification tasks, they have notable challenges. ViT models, for instance, typically require large datasets for effective training, as they do not perform well on small datasets compared to CNNs. This is primarily due to the large number of parameters in Transformers, which require extensive training data to generalize effectively. Additionally, Transformers, while excellent at capturing global information, often struggle with local feature extraction, which is crucial for tasks such as fine-grained art style identification.
@@ 2.3 Related Research on CNN-Transformer Fusion
One prominent approach is proposed by Xie et al., who introduced a hybrid model [7]that first uses CNNs to extract low-level features from images and then passes these features to a Transformer for global modeling. This model has demonstrated good performance in various image classification tasks, including facial identification and general object classification. By combining CNNs’ local feature extraction with Transformers’ ability to capture long-range dependencies, the model achieves enhanced performance over using either architecture in isolation.
Liu et al. (2021) also explored the fusion of CNNs and Transformers, focusing on a multi-scale feature fusion approach. In this method, CNNs extract features at different scales, and these features are subsequently fused with the global features captured by the Transformer. This multi-scale approach is particularly useful when dealing with small datasets, as it allows the model to capture both local and global features from images at different resolutions. This fusion technique has shown significant improvements in small-data environments, making it highly relevant for art identification tasks, where high-quality labeled datasets may be scarce.
While the fusion of CNNs and Transformers has shown promising results in various applications, the integration of these models for the specific task of art author classification remains an area in need of further exploration. Existing research on hybrid models primarily focuses on generic image classification tasks, and there has been limited investigation into optimizing CNN-Transformer fusion architectures for the unique challenges posed by artwork identification, such as capturing fine brushstrokes, identifying complex color gradients, and distinguishing between different artistic authors. Therefore, further work is needed to refine the fusion strategies and adapt them to the specific nuances of art identification tasks.
@@ 3.2 Model Architecture
The model proposed in this study is built on a hybrid architecture combining Convolutional Neural Networks (CNNs) and Transformers, designed to leverage the complementary strengths of these two deep learning approaches. The hybrid model aims to improve classification accuracy and robustness, particularly for the nuanced and diverse nature of art identification tasks.
@@
Fusion Strategy: The hybrid architecture employs a cascading fusion approach to combine CNN and Transformer components. After the CNN extracts local features, these features are passed to the Transformer for further processing. The Transformer applies multiple layers of self-attention to integrate and enhance the global dependencies between these local features. Finally, the local and global features are concatenated and passed to a fully connected layer for classification. This fusion strategy allows the model to effectively combine the strengths of both CNNs and Transformers—local feature extraction and global context modeling—resulting in better performance in classifying art authors.
@@ 3.3 Loss Function and Optimization
Loss Function: The model employs the cross-entropy loss function, which is well-suited for multi-class classification tasks. Given the nature of the art identification task, the model outputs a probability distribution for each sample, representing the likelihood of it belonging to a particular author. The cross-entropy loss function calculates the difference between the predicted probability distribution and the true label distribution, guiding the model to minimize this discrepancy during training. To address the class imbalance in the dataset, where some classes (e.g., rare authors) may have fewer samples, we use a weighted cross-entropy loss. The weights are calculated based on the inverse frequency of each class in the dataset, ensuring that the model gives appropriate attention to minority classes.
@@ 3.4 Data Augmentation and Regularization
Regularization: To further combat overfitting and improve the model’s generalization to unseen data, we employ dropout regularization during training, with a keep probability of 0.5. This technique randomly deactivates certain neurons during training, preventing the model from becoming overly reliant on any single feature and promoting more robust learning. Additionally, we apply L2 regularization (also known as weight decay) to the fully connected layers, with a penalty coefficient of 0.01. This further discourages overly large weights, helping the model maintain generalization across different art authors.
@@ 4.3 Experimental Results
The experimental results on both the Chinese and oil painting datasets are summarized in the table 1. From these results, it is clear that the fusion model significantly outperforms both the individual CNN and Transformer models across all evaluation metrics. For the Chinese painting dataset, the fusion model achieved an impressive 9.7% increase in accuracy and a 0.09 increase in F1-score compared to the CNN-only model. Similarly, for the oil painting dataset, the fusion model outperformed the CNN and Transformer models by 7.1% in accuracy and 0.05 in F1-score. These results highlight the importance of combining the strengths of both CNNs and Transformers for art identification tasks, where both local and global features are essential for accurate classification.
@@ 4.4 Experimental Analysis
Fusion Model (CNN + Transformer): The fusion model, which combines the local feature extraction capability of the CNN with the global modeling power of the Transformer, achieves the best overall performance on both datasets. By integrating the strengths of both architectures, the fusion model captures both local details and global context effectively. This allows the model to handle the nuances of different art authors more effectively. The significant improvements in accuracy and F1-score demonstrate the effectiveness of this approach, particularly when dealing with complex art forms like Chinese and oil paintings.
Furthermore, the fusion model exhibits improved robustness, especially in small-data environments. The integration of both local and global features helps the model generalize better, even when fewer training examples are available, which is particularly important in art identification tasks where labeled data can be limited.
@@ 4.5 Discussion
The fusion of CNN and Transformer allows the model to fully leverage the strengths of both. CNNs are adept at extracting local details from images, such as brushstrokes and color distribution, while Transformers can capture global dependencies and understand the overall artistic author intent. Through the cascading fusion strategy, the model can focus on both local details and global context, thereby improving classification accuracy.
@@ 5 Conclusion
This study proposes an innovative fusion model that combines Convolutional Neural Networks (CNN) and Transformers to address the challenges of art author identification. CNNs, known for their exceptional ability to capture local features such as textures, brushstrokes, and color distributions, are used as the initial step to process the artwork images. Transformers, on the other hand, excel at modeling long-range dependencies and global relationships, which are crucial for understanding the overall composition and style of an artwork. By integrating these two powerful models, the proposed fusion approach leverages the strengths of both CNNs and Transformers, allowing the model to effectively extract detailed local features while also capturing the broader contextual information.
@@
Additionally, the proposed model is especially effective in small-data environments, where the ability to generalize from limited data is crucial. The fusion model not only provides a significant boost in performance but also shows potential in enhancing the identification of subtle artistic nuances that are often challenging for standalone models.
Looking ahead, future research will focus on optimizing the model architecture to further improve its accuracy and efficiency. This could involve exploring more advanced convolutional and Transformer architectures, such as deeper CNN networks or novel Transformer variants. Additionally, the incorporation of multimodal information, including textual descriptions or historical data about artists, could further enrich the model’s understanding of artworks, providing a more holistic approach to art identification. Expanding the dataset to include a wider variety of art authors will also be a key area for future work, helping to improve the model’s generalizability and adaptability across different domains of art.